45.2g of Iron metal reacts with excess Oxygen gas to produce 58.1g of



Fe304



a. What is the theoretical yield of Fe304? b. What is the percent yield of this reaction?

1
Expert's answer
2022-05-13T05:20:04-0400

Solution:

The molar mass of iron (Fe) is 55.845 g/mol

Therefore,

Moles of Fe = (45.2 g Fe) × (1 mol Fe / 55.845 g Fe) = 0.8094 mol Fe


Balanced chemical equation:

3Fe(s) + 2O2(g) → Fe3O4(s)

According to stoichiometry:

3 mol of Fe produce 1 mol of Fe3O4

Thus, 0.8094 mol of Fe produce:

(0.8094 mol Fe) × (1 mol Fe3O4 / 3 mol Fe) = 0.2698 mol Fe3O4


The molar mass of Fe3O4 is 231.533 g/mol

Therefore,

Mass of Fe3O4 = (0.2698 mol Fe3O4) × (231.533 g Fe3O4 / 1 mol Fe3O4) = 62.4676 g Fe3O4

Mass of Fe3O4 = 62.4676 g = 62.5 g

The theoretical yield of Fe3O4 is 62.5 grams


Percent yield = (Actual yield / Theoretical yield) × 100%

Percent yield = (58.1 g / 62.5 g) × 100% = 92.96%

The percent yield of this reaction is 92.96%


Answers:

(a): The theoretical yield of Fe3O4 is 62.5 grams

(b): The percent yield of this reaction is 92.96%